Sam farms a 400-acre farm. Two hundred acres of the land are particularly rich in a nutrient that is favorable to growing corn but has no benefits for wheat. Sam can grow 30 bushels of corn or 25 bushels of wheat on each of these special acres. The other 200 acres are not rich in the nutrient, so Sam can grow only 10 bushels of corn or the same 25 bushels of wheat per acre.

Suppose that Sam considers starting with all of his land in corn. Then he thinks about increasing the amount of wheat that he grows. What is the opportunity cost, in bushels of corn, of the first bushel of wheat that he grows?
1
Expert's answer
2012-05-31T08:14:30-0400
The opportunity cost is the cost of an alternative that must be forgone in order to pursue a certain action. In the other words, the benefits someone could have received by
taking an alternative action.
But first we need to calculate the bushels of corn as Sam decided to grow the corn on his land first.
200 × 30 = 6000 bushels of corn (he can get from the land rich in a nutrient)
200 × 10 = 2000 bushels of corn (he can get from the land not rich in a nutrient)
6000 + 2000 = 8000 bushels of corn (altogether)
On the other hand, if he decided to grow the wheat he would get:
200 × 25 × 2 = 10000 bushels of wheat (as from both part of land he can get the same 25 bushels of wheat).
We need to compare the opportunity cost of one bushel of corn and one bushel of wheat to be able to know which choice
between two options must be made. The risk that you could achieve greater
"benefits" (be they monetary or otherwise) with another option is the
opportunity cost, so we need to know if growing of the wheat will be better
choice that the growing of corn:
8000 ÷ (30+10) = 200
10000 ÷ (25+25) = 200
The opportunity cost is the difference in return between a chosen option and alternative one. In our example we see that
the opportunity cost will be 0, so in the given case the both options will give
the same result / return for the gardener.